OpenStudy (anonymous):

Happy Beans coffee shop decides to blend two coffees: a) Turkish, at $11.00 per pound, and b) Brazilian, at $14.00 per pound. The shop plans to make a blend that sells for $12.00 per pound. How much of each to make 50 pounds of this blend?

OpenStudy (tkhunny):

First, name things: T = # of Turkish B = # of Brazillian We know T + B = 50 <== Equation for Pounds We know T(11) + B(14) = 50(12) <== Equation for Price Now, we just need a way to solve for T and B. Suggestions?

OpenStudy (anonymous):

why not find a way to get 6 dollars of each blend?

OpenStudy (anonymous):

then add both of those together?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

Thanks everyone. Between all the answers from everyone i was able to figure out the problem.